OpenStudy (anonymous):

Simplify the expression: (3xy)^-1

jimthompson5910 (jim_thompson5910):

Similar problem: \[\Large (7xyz)^{-1} = \frac{1}{(7xyz)^1}\] \[\Large (7xyz)^{-1} = \frac{1}{7xyz}\]

OpenStudy (anonymous):

A. - 1 3xy B. -3xy C. 1 3xy D. 3xy

jimthompson5910 (jim_thompson5910):

so which one do you think it is based off that example i gave you?

OpenStudy (anonymous):

c

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

assuming C is \[\Large \frac{1}{3xy}\]